OpenStudy (anonymous):

3) How long does it take the apple to reach your friend? (**HINT: The time would be the x-coordinate of the highest point.) Double click on the space below to show your work. h(t) = – 16t2 + 38.4t + 0.96 H= Height. T= Time (seconds). I WILL FAN AND MEDAL FOR ANSWERS AND WORK.

OpenStudy (anonymous):

Set the equation = 0. Then use quadratic formula.

OpenStudy (anonymous):

It is supposed to be h(t) = -16t^2 + 38.4t + 0.96

OpenStudy (anonymous):

Ok! Set the equation = 0.

OpenStudy (anonymous):

\(\huge -16t^2+38.4t+0.96=0\)

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

Then it would be \[-38\pm \sqrt{38.4^{2}-4(-16)(0.96)} / 2(-16)\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

The beginning is supposed to be -38.4 not just -38.

OpenStudy (anonymous):

Yes! Now solve for t. :) And since it is time, your answer shouldn't be the negative one.

OpenStudy (anonymous):

Let me know if you got an answer, so I may check if it's right. :)

OpenStudy (anonymous):

So then I get: \[-38.4\pm \sqrt{1,536} / -32\] I'm confused on what to do after this step.

OpenStudy (anonymous):

\(\huge \frac{-38.4±\sqrt {1536}}{-32}\) Frind the square root of 1536 first.

OpenStudy (anonymous):

\[16\sqrt{6}\]

OpenStudy (anonymous):

\(\huge \frac{-38.4±16\sqrt {6}}{-32}\rightarrow\) \(\huge \frac{-38.4+16\sqrt {6}}{-32} \) & \(\huge \frac{-38.4-16\sqrt {6}}{-32} \)

OpenStudy (anonymous):

Should it be in radical form or could it be in decimal form?

OpenStudy (anonymous):

So how do we get the answer from there? It should be in decimal form since we are trying to find how many seconds it takes the apple to reach your friend.

OpenStudy (anonymous):

Great! Now get a calculator and solve for it. Notice that \(\sqrt{6} =2.45\).

OpenStudy (anonymous):

\(\huge \frac{-38.4+16(2.45)}{-32}\) &\(\huge \frac{-38.4-16(2.45)}{-32}\)

OpenStudy (anonymous):

Still confusing?

OpenStudy (anonymous):

I got -0.02

OpenStudy (anonymous):

Or should it be positive 0.02? @mathway

OpenStudy (anonymous):

You're supposed to get \(\huge \color{red}{ \text{TWO}}\) solutions. and \(-0.02\) is one. And since it is negative, it can't be the time. So solve for the other one.

OpenStudy (anonymous):

2.4 seconds rounded.

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

No problem! :)